OpenStudy (anonymous):

2x/x^2-4 = 4/x^2-4 - 3/x+2

OpenStudy (valpey):

\[x^2-4=(x-2)(x+2)\]Try to get each term with the same denominator (or none at all). Make a note that x cannot equal 2 or -2 because the equation is not defined in those cases.

OpenStudy (anonymous):

can you explain it please?

OpenStudy (anonymous):

@Valpey can you help me work this problem

OpenStudy (valpey):

\[\frac{2x}{x^2-4}=\frac{4}{x^2-4}-\frac{3}{x+2}\]But as I was showing: \[x^2-4=(x-2)(x+2)\]So then we have:\[\frac{2x}{(x-2)(x+2)}=\frac{4}{(x-2)(x+2)}-\frac{3}{x+2}\]But I want to cancel out all those denominators so I....?

OpenStudy (anonymous):

multiply both sides by (x-2)(x+2)?

OpenStudy (valpey):

Good job! What would we have then?

OpenStudy (anonymous):

2x=4-3

OpenStudy (valpey):

Not quite... Re-do the third term.

OpenStudy (anonymous):

2x=4-3(x-2)?

OpenStudy (valpey):

Yes!

OpenStudy (anonymous):

thank you!!!